# Universal control of four singlet-triplet qubits

### v2

@@ Abstract
The coherent control of interacting spins in semiconductor quantum dots is of strong interest for quantum information processing as well as for studying quantum magnetism from the bottom up. On paper, individual spin-spin couplings can be independently controlled through gate voltages, but nonlinearities and crosstalk introduce significant complexity that has slowed down progress in past years. Here, we present a $2\times 4$ germanium quantum dot array with full and controllable interactions between nearest-neighbor spins. As a demonstration of the level of control, we define four singlet-triplet qubits in this system and show two-axis single-qubit control of all qubits and SWAP-style two-qubit gates between all neighbouring qubit pairs. Combining these operations, we experimentally implement a circuit designed to generate and distribute entanglement across the array. These results highlight the potential of singlet-triplet qubits as a competing platform for quantum computing and indicate that scaling up the control of quantum dot spins in extended bilinear arrays can be feasible.
@@
The coherent control of a large-scale array of spins in the solid state represents a major challenge in the field of quantum-coherent nanoscience [1, 2, 3, 4]. As a quintessential platform for quantum spin control, the lithographically-defined semiconductor quantum dot has shown great promise both for fault-tolerant digital quantum computationo [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and for analog quantum simulation of emergent quantum phenomena [16, 17, 18, 19]. Nevertheless, the inherent nanoscale dimensions of the devices, the geometrical constraints in integrating all the required components, and the necessity of employing high-frequency electromagnetic fields in cryogenic environments present important challenges for the integration and control a large number of spins.
@@
Already, significant efforts have been undertaken to tackle these challenges. For single-spin qubits, the number of coherently controlled spins has been scaled up to six in a one-dimensional array [11] and four in a two-dimensional array [10]. A six-dot linear array was also used to achieve universal control of two qubits that are each encoded in a subspace of three electron spins distributed over three dots [15]. For singlet-triplet qubits defined in a subspace of two spins across two dots, recent progress includes the individual control of three to four uncoupled qubits [20, 21] and the operation of a single qubit in a $3\times 3$ quantum dot array [22].
@@
Similar to exchange-only qubits, singlet-triplet qubits [6, 23, 24, 25, 13, 26, 14, 27, 28, 29, 30, 21] allow fully electrical qubit control using baseband voltage pulses. The use of baseband-only control signals can avoid commonly encountered problems of single-spin qubits such as microwave heating effects [31, 11, 32] and may furthermore alleviate crosstalk effects [33]. Singlet-triplet qubits also map naturally to the spin-readout basis in Pauli spin blockade (PSB), which is a common method for spin-to-charge conversion in quantum dots [34, 24]. By using pulse optimization, single-qubit control fidelities of single-triplet qubits have exceeded $99\%$ [28], whereas two-qubit gate fidelities relying on the relatively weak capacitive (Coulomb) interaction reached 72-90% [13, 14]. In theory, the two-qubit gate fidelity can be further improved by replacing the capacitive coupling with the stronger exchange coupling [29], although it has been little investigated in experiments [35]. Despite this progress, universal control of more than two interacting singlet-triplet qubits remains yet to be achieved. Recently, a controlled number of charge carriers were loaded in $2\times 4$ arrays, a $4\times 4$ array and a $1\times 12$ array [36, 37, 38, 39, 40]. These advances set the stage for exploring the operation of three or more interacting singlet-triplet qubits experimentally.
@@
Figure 1: Device and energy spectroscopy. a, Schematic drawing showing the Ge/SiGe heterostructure and three layers of gate electrodes on top to define the quantum dot ladder and sensing dots: screening gates (purple), plunger gates (red), barrier gates (green). Ohmic contacts (light green) extend towards the Ge quantum well in which the holes are confined. The aluminum oxide dielectric between different gate layers is omitted for clarity. b, False-colored scanning electron microscope image of a device nominally identical to that used in the measurements, where the in-plane position of the 8 quantum dots is indicated with numbers 1-8 in circles. Charge sensors close to the ladder corners are labeled within larger circles. The plunger (red) and barrier (green) gates of the quantum dots are labeled outside the image. A schematic of the ladder structure of the quantum dots is shown on top, with Q1-Q4 formed by vertical double quantum dots (DQD). c-e, The energy levels of two-spin states in a DQD as a function of energy detuning $\varepsilon_{ij}$ between dot $i$ and $j$ with from left to right the case $J(\varepsilon_{ij}=0)<\overline{E}_{z}$, $J(\varepsilon_{ij}=0)=\overline{E}_{z}$, $J(\varepsilon_{ij}=0)>\overline{E}_{z}$. The dashed black circles denote the positions of $S-T_{-}$ anticrossings. f-i, The measured energy spectra that probe the positions of the $S-T_{-}$ anticrossings as a function of the detuning and the barrier gate voltage for each vertical DQD. The color scale shows the measured spin triplet probability $P_{T}$ after initializing a vertical DQD in a singlet state (in (0,2) or (2,0)) and applying a gate voltage pulse (20 ns ramp in, 40 ns wait time, 0 ns ramp out) to the detuning shown on the horizontal axis, for different $\text{v}B_{ij}$.
@@ I Germanium quantum dot ladder
As shown in Fig. 1a and b, the 2x4 quantum dot ladder is fabricated in a germanium quantum-well heterostructure [50]. The gate pattern and substrate have the same design as that in ref. [38]. The eight quantum dots are labeled with numbers 1 to 8 and the four charge sensors to measure the charge states in the quantum dots are labeled $\text{S}_{\text{TL}}$, $\text{S}_{\text{TR}}$, $\text{S}_{\text{BL}}$ and $\text{S}_{\text{BR}}$, respectively. The quantum dot potentials are controlled by plunger gates $\text{P}_{i}$, and the interdot or dot-sensor tunnel couplings are controlled by barrier gates $\text{B}_{ij}$ or $\text{B}_{i}$, with $i$ or $j$ denoting the corresponding quantum dot number. Linear combinations of plunger gate voltages $\{P_{i}\}$ allow us to set the overall electrochemical potential of each DQD $\mu_{ij}=(\text{v}P_{i}+\text{v}P_{j})/2$ and the interdot detuning $\varepsilon_{ij}=(\text{v}P_{i}-\text{v}P_{j})/2$. The prefix “v” indicates that the physical gate voltages are virtualized to compensate the crosstalk on the dot potentials [16] (see Supplementary Information section III for the virtual gate matrix). Single-hole occupation of each quantum dot in the array is confirmed by measuring the charge stability diagrams using sensors $\text{S}_{\text{BL}}$ and $\text{S}_{\text{BR}}$ (see Supplementary Information section II). All plunger and interdot barrier gates are connected to a bias tee to allow both DC voltages and voltage pulses to be applied.
@@ II Singlet-triplet qubit and energy spectroscopy
Figure 2: Universal single-qubit control of four singlet-triplet qubits. a,b, Illustration of $x$-axis and $z$-axis evolution of $S-T_{-}$ qubits in the Bloch sphere. The straight blue and orange arrows show the corresponding rotation axis. The $x$-axis rotations are set by the $S-T_{-}$ coupling, $\Delta_{ST_{-}}$. For large $J$ such that $J-\overline{E}_{z}\gg\Delta_{ST_{-}}$, the rotation axis tilts towards the $z$-axis. c-f, Experimental results for $x$-axis rotations of each qubit, showing measured triplet probabilities $P_{T}$ as a function of $t_{\text{wait}}$ and the detuning voltage $\varepsilon_{ij}$. g-j, Experimental results for $z$-axis rotations of each qubit, showing $P_{T}$ as a function of $t_{\text{wait}}$ and the barrier voltage change $\delta\text{v}B_{ij}$. k,l, The pulse schemes used for $x$-axis control (k) and $z$-axis control (l). In experiments, the detuning pulse in k and l has a 20 ns ramp (not shown) from (2,0) to (1,1) to ensure adiabaticity with respect to the $S$-$T_{0}$ splitting. m, The rotation frequency $f_{ST_{-}}$ of each qubit as a function of the magnetic field strength $B$. The data are acquired using the fast Fourier transform (FFT) of time-domain oscillations (initiated using a ramp time of 100 ns, see Supplementary Information section V for more details and for the time-domain data of m). n, The rotation frequency $f_{ST_{-}}$ of each qubit as a function of $\delta\text{v}B_{ij}$, extracted from panels g-j using FFT. The data points for Q2 and Q3 show some steps as the barrier gate voltage is changed, which are limited by the FFT precision. The solid lines are fits of the data with an exponential function. o-p, The measured triplet probabilities $P_{T}$ of long-time evolutions around the $x$- (o) and $z$-axis (p) for Q1-Q4. The data is vertically shifted for clarity. In o, the detuning is around 0 mV. In p, we take the data with the same detuning but with different barrier voltage pulses: $\delta\text{v}B_{15}=-20$ mV, $\delta\text{v}B_{26}=-80$ mV, $\delta\text{v}B_{37}=-80$ mV, and $\delta\text{v}B_{48}=-30$ mV. The barrier voltages are chosen to achieve the condition $J-\overline{E}_{z}\gg\Delta_{ST_{-}}$ depending on the gate tunability. For more measurement details, see Supplementary Information section II.
We encode the qubit into the two-spin singlet-triplet states, |$\ket{S}$ and |$\ket{T_{-}}$, of the DQDs along the rungs of the quantum dot ladder, with the singlet |$\ket{S}=(\ket{\uparrow\downarrow}-\ket{\downarrow\uparrow})/\sqrt{2}$ and the lowest-energy triplet |$\ket{T_{-}}=\ket{\downarrow\downarrow}$. Thus Q1, Q2, Q3 and Q4 are formed using DQD 1-5, 2-6, 3-7 and 4-8. Qubit readout is achieved by pulsing the corresponding DQD to the PSB regime. This regime converts the singlet and triplet states into distinct charge states, which are then measured through the charge sensor. The single-qubit Hamiltonian can be written as
@@
where $\sigma_{x}$ and $\sigma_{z}$ are the Pauli matrices, $J=J(\varepsilon_{ij},\text{v}B_{ij})$ is a function of both the detuning $\varepsilon_{ij}$ and the barrier gate voltage $\text{v}B_{ij}$, and $\overline{E}_{z}=\overline{g}_{ij}\mu_{B}B$ is the average Zeeman energy of the two hole spins in the DQD, with $\overline{g}_{ij}$ the average $g$-factor, $\mu_{B}$ the Bohr magneton, and $B$ the magnetic field strength. Unless indicated otherwise, an in-plane magnetic field (up to alignment precision) of $B=10$ mT is applied to the device. The intrinsic spin-orbit interaction for holes in germanium couples the states |$\ket{S}$ and |$\ket{T_{-}}$ with an energy $\Delta_{ST_{-}}$.
@@
Experimentally, we probe the position of the avoided crossings as follows. First, we initialize one of the qubits to a singlet by diabatically pulsing from (0,2) or (2,0) to the detuning $\varepsilon_{ij}$ in (1,1). After waiting for a certain time, we pulse the qubit back to the PSB regime to record the triplet probability. When the pulse takes the system to an anticrossing, the singlet will evolve into a triplet during the waiting time (a 40 ns duration is chosen to obtain a sizable triplet probability). Performing such measurements as a function of the barrier gate voltage $\text{v}B_{ij}$ that controls $J$ for each qubit, results in the parabola-like patterns, also called spin mixing maps [51, 22], in Fig. 1f-i. As expected, when $\text{v}B_{ij}$ is tuned from positive to negative, $J$ increases and the positions of the $S-T_{-}$ anticrossings move inwards before disappearing. The asymmetry visible in panel g in particular can arise from imperfect virtualization of the barrier gates or from a detuning-dependent Zeeman energy [52] (see Supplementary Information section IV).
@@ III Universal single-qubit control
With the knowledge of the energy spectrum of the four $S-T_{-}$ qubits, we next implement the two-axis control of each qubit, which is necessary and sufficient for universal single-qubit control. By operating the qubit in the regime where $J=\overline{E}_{z}$, the first term of Eq. 1 goes to zero and $\Delta_{ST_{-}}$ rotates the qubit around the $x$-axis in the Bloch sphere, as shown in Fig. 2a. Furthermore, we tune the barrier voltage to obtain $J=\overline{E}_{z}$ at zero detuning, which is a symmetry point where the effect of detuning noise is strongly suppressed [53, 54]. The pulse scheme for testing $x$-axis control is shown in Fig. 2k: first we initialize the qubit into a singlet by starting in the (2,0) (or (0,2)) regime, then pulse the detuning to the center of the (1,1) regime where $J(\varepsilon_{ij}=0)=\overline{E}_{z}$, next allow the qubit to evolve for a variable time $t_{\text{wait}}$, and finally pulse the detuning back to a point in the (2,0) (or (0,2)) regime for spin readout via PSB. The measured rotations of Q1-Q4 as a function of the pulsed detuning position are shown in Fig. 2c-f. With proper calibration, the qubits rotate around the $x$-axis at zero detuning (this condition occurs slightly away from the point $\varepsilon_{ij}=0$ extracted from charge stability diagrams).
To realize $z$-axis control, we operate the Hamiltonian in the opposite limit where we increase $J$ to such a large value that the effect of spin-orbit coupling becomes negligible. As a result, the qubit is rotated around an axis close to the $z$-axis, with a frequency proportional to $J-\overline{E}_{z}$. The rotation is never exactly around the $z$-axis due to the presence of a finite $\Delta_{ST_{-}}$, yet, sufficiently orthogonal control is possible when $(J-\overline{E}_{z})\gg\Delta_{ST_{-}}$ (furthermore, by finely tuning the magnet field direction, $\Delta_{ST_{-}}=0$ can be achievable [49, 46]). In experiments, we perform a Ramsey-like pulse sequence to demonstrate $z$-axis control. As illustrated in Fig. 2b and l, we first initialize the qubit into a singlet, perform a $\pi/2$ rotation around the $x$-axis of duration $t_{\pi/2}$, and then change $J$ diabatically by pulsing the corresponding barrier gate by an amount $\delta\text{v}B_{ij}$ to implement a $z$-axis rotation. Throughout the barrier gate pulse, we aim to stay at the symmetry point to minimize the sensitivity to charge noise [53, 54]. Finally, we perform another $\pi/2$ operation around the $x$-axis and project the qubit into the $S-T_{-}$ basis for spin readout.

### v3

@@ Abstract
The coherent control of interacting spins in semiconductor quantum dots is of strong interest for quantum information processing as well as for studying quantum magnetism from the bottom up. Here, we present a $2\times 4$ germanium quantum dot array with full and controllable interactions between nearest-neighbor spins. As a demonstration of the level of control, we define four singlet-triplet qubits in this system and show two-axis single-qubit control of each qubit and SWAP-style two-qubit gates between all neighbouring qubit pairs, yielding average single-qubit gate fidelities of 99.49(8)-99.84(1)and Bell state fidelities of 73(1)-90(1)%. Combining these operations, we experimentally implement a circuit designed to generate and distribute entanglement across the array. A remote Bell state with a fidelity of 75(2)and concurrence of 22(4)is achieved. These results highlight the potential of singlet-triplet qubits as a competing platform for quantum computing and indicate that scaling up the control of quantum dot spins in extended bilinear arrays can be feasible.
@@
The coherent control of a large-scale array of spins in the solid state represents a major challenge in the field of quantum-coherent nanoscience [1, 2, 3, 4]. As a quintessential platform for quantum spin control, the lithographically-defined semiconductor quantum dot has shown great promise both for fault-tolerant digital quantum computation [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15] and for analog quantum simulation of emergent quantum phenomena [16, 17, 18, 19]. Nevertheless, the inherent nanoscale dimensions of the devices, the geometrical constraints in integrating all the required components, and the necessity of employing high-frequency electromagnetic fields in cryogenic environments present important challenges for the integration and control of a large number of spins.
@@
Already, significant efforts have been undertaken to tackle these challenges. For single-spin qubits, the number of coherently controlled interacting spins has been scaled up to six in a one-dimensional array [11] and four in a two-dimensional array [10]. A six-dot linear array was also used to achieve universal control of two qubits that are each encoded in a subspace of three electron spins distributed over three dots [15]. For singlet-triplet qubits defined in a subspace of two spins across two dots, recent progress includes the individual control of three to four uncoupled qubits [20, 21] and the operation of a single qubit in a $3\times 3$ quantum dot array [22].
@@
Similar to exchange-only qubits, singlet-triplet qubits [6, 23, 24, 25, 13, 26, 14, 27, 28, 29, 30, 21] allow fully electrical qubit control using baseband voltage pulses. The use of baseband-only control signals can avoid commonly encountered problems of single-spin qubits such as microwave heating effects [31, 11, 32] and may furthermore alleviate crosstalk effects [33]. Singlet-triplet qubits also map naturally to the spin-readout basis in Pauli spin blockade (PSB), which is a common method for spin-to-charge conversion in quantum dots [34, 24]. By using pulse optimization, single-qubit control fidelities of singlet-triplet qubits have exceeded $99\%$ [28], whereas two-qubit gate fidelities relying on the relatively weak capacitive (Coulomb) interaction reached 72-90% [13, 14]. In theory, the two-qubit gate fidelity can be further improved by replacing the capacitive coupling with the stronger exchange coupling [29], although it has been little investigated in experiments [35]. Despite this progress, universal control of more than two interacting singlet-triplet qubits remains yet to be achieved. Recently, a controlled number of charge carriers were loaded in $2\times 4$ arrays, a $4\times 4$ array and a $1\times 12$ array [36, 37, 38, 39, 40]. These advances set the stage for exploring the operation of three or more interacting singlet-triplet qubits experimentally.
@@
Figure 1: Device and energy spectroscopy. a, Schematic drawing showing the Ge/SiGe heterostructure and three layers of gate electrodes on top to define the quantum dot ladder and sensing dots: screening gates (purple), plunger gates (red), barrier gates (green). Ohmic contacts (gray) extend towards the Ge quantum well in which the holes are confined. The aluminum oxide dielectric between different gate layers is omitted for clarity. b, False-colored scanning electron microscope image of a device nominally identical to that used in the measurements, where the in-plane position of the 8 quantum dots is indicated with numbers 1-8 in circles. Charge sensors close to the ladder corners are labeled within larger circles. The plunger (red) and barrier (green) gates of the quantum dots are labeled outside the image. A schematic of the ladder structure of the quantum dots is shown on top, with Q1-Q4 formed by vertical double quantum dots (DQD). c-e, The energy levels of two-spin states in a DQD as a function of energy detuning $\varepsilon_{ij}$ between dot $i$ and $j$ with from left to right the case $J(\varepsilon_{ij}=0)<\overline{E}_{z}$, $J(\varepsilon_{ij}=0)=\overline{E}_{z}$, $J(\varepsilon_{ij}=0)>\overline{E}_{z}$. The dashed black circles denote the positions of $S-T_{-}$ anticrossings. f-i, The measured energy spectra that probe the positions of the $S-T_{-}$ anticrossings as a function of the detuning and the barrier gate voltage for each vertical DQD at $B=5$ mT. The color scale shows the measured spin triplet probability $P_{T}$ after initializing a vertical DQD in a singlet state (in (0,2) or (2,0)) and applying a gate voltage pulse (20 ns ramp in, 40 ns wait time, 0 ns ramp out) to the detuning shown on the horizontal axis, for different $\text{v}b_{ij}$. The cartoons on top of panels f-i represent the eight dots, and the dark grey line indicates which exchange coupling is active in the panel below.
@@ I Germanium quantum dot ladder
As shown in Fig. 1a and b, the 2x4 quantum dot ladder is fabricated in a germanium quantum-well heterostructure [50]. The gate pattern and substrate have the same design as that in ref. [38]. The eight quantum dots are labeled with numbers 1 to 8 and the four charge sensors to measure the charge states in the quantum dots are labeled $\text{S}_{\text{TL}}$, $\text{S}_{\text{TR}}$, $\text{S}_{\text{BL}}$ and $\text{S}_{\text{BR}}$, respectively. The quantum dot potentials are controlled by plunger gates $\text{P}_{i}$, and the interdot or dot-sensor tunnel couplings are controlled by barrier gates $\text{b}_{ij}$ or $\text{b}_{i}$, with $i$ or $j$ denoting the corresponding quantum dot number. Linear combinations of plunger gate voltages $\{P_{i}\}$ allow us to set the overall electrochemical potential of each DQD $\mu_{ij}=(\text{v}P_{i}+\text{v}P_{j})/2$ and the interdot detuning $\varepsilon_{ij}=(\text{v}P_{i}-\text{v}P_{j})/2$. The prefix “v” indicates that the physical gate voltages are virtualized to compensate the crosstalk on the dot potentials [16] (see Supplementary Information section II for the virtual gate matrix). Single-hole occupation of each quantum dot in the array is confirmed by measuring the charge stability diagrams using sensors $\text{S}_{\text{BL}}$ and $\text{S}_{\text{BR}}$ (see Extended Data Fig. 1). All plunger and interdot barrier gates are connected to a bias tee to allow both DC voltages and voltage pulses to be applied.
@@ II Singlet-triplet qubit and energy spectroscopy
We encode the qubit into the two-spin singlet-triplet states, |$\ket{S}$ and |$\ket{T_{-}}$, of the DQDs along the rungs of the quantum dot ladder, with the singlet |$\ket{S}=(\ket{\uparrow\downarrow}-\ket{\downarrow\uparrow})/\sqrt{2}$ and the lowest-energy triplet |$\ket{T_{-}}=\ket{\downarrow\downarrow}$. Thus Q1, Q2, Q3 and Q4 are formed using DQD 1-5, 2-6, 3-7 and 4-8. Qubit readout is achieved by pulsing the corresponding DQD to the PSB regime, i.e. in the (0,2) or (2,0) regime but close to (1,1). This regime converts the singlet and triplet states into distinct charge states, which are then measured through the charge sensor (see Extended Data Fig. 1 for details). The single-qubit Hamiltonian can be written as
@@
where $\sigma_{x}$ and $\sigma_{z}$ are the Pauli matrices, $J=J(\varepsilon_{ij},\text{v}b_{ij})$ is a function of both the detuning $\varepsilon_{ij}$ and the barrier gate voltage $\text{v}b_{ij}$, and $\overline{E}_{z}=\overline{g}_{ij}\mu_{B}B$ is the average Zeeman energy of the two hole spins in the DQD, with $\overline{g}_{ij}$ the average $g$-factor, $\mu_{B}$ the Bohr magneton, and $B$ the magnetic field strength. Unless indicated otherwise, an in-plane magnetic field (up to alignment precision) of $B=5$ or $10$ mT is applied to the device. The intrinsic spin-orbit interaction for holes in germanium couples the states |$\ket{S}$ and |$\ket{T_{-}}$ with an energy $\Delta_{ST_{-}}$.
@@
Experimentally, we probe the position of the avoided crossings as follows. First, we initialize one of the qubits to a singlet by pulsing from (0,2) or (2,0) to the detuning $\varepsilon_{ij}$ in (1,1). A 20 ns ramp-in time is used to ensure adiabaticity with respect to the tunnel coupling, which is around 2 GHz, and diabaticity with respect to the $S-T_{-}$ anticrossing. After waiting for a certain time, we pulse the qubit back to the PSB regime to record the triplet probability (see Extended Figure 1 for full details). When the pulse takes the system to an anticrossing, the singlet will evolve into a triplet during the waiting time (of 40 ns duration, close to a $\pi$ rotation is chosen to obtain a sizable triplet probability). Performing such measurements as a function of the barrier gate voltage $\text{v}b_{ij}$ that controls $J$ for each qubit, results in the parabola-like patterns, also called spin mixing maps [51, 22], in Fig. 1f-i. As expected, when $\text{v}b_{ij}$ is tuned from positive to negative, $J$ increases and the positions of the $S-T_{-}$ anticrossings move inwards before disappearing. The asymmetry visible in these panels can arise from imperfect virtualization of the barrier gates or from a detuning-dependent Zeeman energy [52] (see Supplementary Information section III).
@@ III Universal single-qubit control
With the knowledge of the energy spectrum of the four $S-T_{-}$ qubits, we next implement the two-axis control of each qubit, which is necessary and sufficient for universal single-qubit control. By operating the qubit in the regime where $J=\overline{E}_{z}$, the first term of Eq. 1 goes to zero and $\Delta_{ST_{-}}$ rotates the qubit around the $x$-axis in the Bloch sphere, as shown in Fig. 2a. Furthermore, we tune the barrier voltage to obtain $J=\overline{E}_{z}$ at zero detuning, which is a symmetry point where the effect of detuning noise is strongly suppressed [53, 54]. The pulse scheme for testing $x$-axis control is shown in Fig. 2k: first we initialize the qubit into a singlet by starting in the (2,0) (or (0,2)) regime, then pulse the detuning to the center of the (1,1) regime where $J(\varepsilon_{ij}=0)=\overline{E}_{z}$, next allow the qubit to evolve for a variable time $t_{\text{wait}}$, and finally pulse the detuning back to a point in the (2,0) (or (0,2)) regime for spin readout via PSB. The measured rotations of Q1-Q4 as a function of the corresponding barrier gate voltage are shown in Fig. 2c-f. By choosing the point where the oscillation speed is the slowest, i.e., at the $S-T_{-}$ anticrossing, the qubits rotate around the $x$-axis. Long-timescale $x$-rotations are shown in Extended Data Fig. 3. At $B=5$ mT, the dephasing times are in the range of 1.5-2.2 $\mu$s for Q1-Q4, mostly limited by low-frequency or quasi-static noise.
To realize $y$-axis control, we use the relationship $\sqrt{Y}=XH$, where $\sqrt{Y}$ stands for a $\pi/2$ rotation around the $y$-axis, $X$ for a $\pi$ rotation around the $x$-axis, and $H$ refers to the Hadamard gate (for direct $z$-axis control, see Extended Data Fig. 2). To obtain a Hadamard gate, we slightly increase $J$ to a value where $(J-\overline{E}_{z})=\Delta_{ST_{-}}$. The qubit then rotates around an axis halfway between the $x$ and $z$ axes. A rotation angle of $\pi$ then corresponds to a Hadamard gate. In order to calibrate the rotation axis and rotation angle of the $H$ gate, we concatenate two $\sqrt{Y}$ gates and evaluate the probability of having flipped the qubit (Fig. 2b). Specifically, we first initialize the qubit into a singlet, then change $J$ diabatically by pulsing the corresponding barrier gate by an amount $\delta\text{v}b_{ij}$ for a time $t_{\text{wait}}$ to implement the $H$ gate, and finally apply a $X$ gate (we aim to stay at the detuning symmetry point when pulsing the barrier to minimize the sensitivity to charge noise [53, 54]). This combination is repeated such that a $Y$ gate is expected for the right choice of $\delta\text{v}b_{ij}$ and $t_{\text{wait}}$. This procedure is illustrated for Q2 in Fig. 2g. The white dot shows the position where the sequence of Fig. 2b produces a $Y$ gate. Running only the first half of the sequence implements a $\sqrt{Y}$ gate. A corresponding numerical simulation result is shown in the inset of Fig. 2g, showing the same pattern as the experimental result.